# Membership Inference Attacks on Recommender System: A Survey

### v1

@@ III-C Evaluation Metric
AUC (ROC AUC). Area under the ROC curve; a threshold-independent summary of performance across all decision thresholds and interpretable as the probability that a randomly chosen positive is scored higher than a randomly chosen negative.
@@
IV Defense
Although research on membership inference attacks (MIAs) for recommender systems has accelerated, effective defenses remain comparatively underdeveloped. For user-level privacy, the Popularity Randomization strategy—randomly recommending from an expanded pool of popular items to non-members—has been shown to reduce attack success and is commonly adopted as a baseline defense [29, 25]. At the interaction level, federated recommenders pose distinct challenges: applying local differential privacy (LDP) [27] can mitigate inference but is often ineffective unless recommendation utility is substantially degraded; instead, Yuan et al.[28] design a simple defense that curbs attack accuracy without sacrificing recommendation performance [28]. For knowledge-graph–based interaction attacks (e.g., MINER), the authors further propose an RGL defense; in addition, standard countermeasures such as noisy embeddings [30] and embedding-based regularizers [28] have been evaluated in related interaction-privacy settings[32]. Differentially private training (DP-SGD) [19, 2] generally provides stronger formal protection, but at a marked cost to model utility. Balancing privacy guarantees with recommendation quality therefore remains a central open problem.
V Future and Direction

### v2

@@ 4.0.3. Metrics
To verify the effectiveness of these MIAs, they are seen as binary classification to be evaluate. We summarize the common evaluation metric. AUC (ROC AUC). Area under the ROC curve; a threshold-independent summary of performance across all decision thresholds and interpretable as the probability that a randomly chosen positive is scored higher than a randomly chosen negative.
@@
5. Defense
Although research on membership inference attacks (MIAs) for recommender systems has accelerated, effective defenses remain comparatively underdeveloped. For user-level privacy, the Popularity Randomization strategy—randomly recommending from an expanded pool of popular items to non-members—has been shown to reduce attack success and is commonly adopted as a baseline defense (Zhang et al., 2021; Wang et al., 2022). At the interaction level, federated recommenders pose distinct challenges: applying local differential privacy (LDP) (Wu et al., 2022) can mitigate inference but is often ineffective unless recommendation utility is substantially degraded; instead, Yuan et al.(Yuan et al., 2023) design a simple defense that curbs attack accuracy without sacrificing recommendation performance (Yuan et al., 2023). For knowledge-graph–based interaction attacks (e.g., MINER), the authors further propose an RGL defense; in addition, standard countermeasures such as noisy embeddings (Zhang et al., 2022) and embedding-based regularizers (Yuan et al., 2023) have been evaluated in related interaction-privacy settings(Zhong et al., 2024). Differentially private training (DP-SGD) (Müllner et al., 2023; Abadi et al., 2016) generally provides stronger formal protection, but at a marked cost to model utility. Balancing privacy guarantees with recommendation quality therefore remains a central open problem.
6. Discussion and Future Direction